A clown visted three shops. He spent $25 at the first shop. At the second shop, he spent half of his remaining money. At the third shop, he spent 1/3 of his remaining money and had $60 left. How much money did the clown have originally?
1
Expert's answer
2010-07-08T04:19:31-0400
Let's consider that clown had x dollars. At the first shop he spent $25. At the second shop he spent half of his remaining money - ((x-25)/2). At the third shop, he spent 1/3 of his remaining money - (x-25-((x-25)/2)/3 and had $60 left. So we can write down an equation and find x:

25+((x-25)/2)+((x-25-((x-25)/2)/3)+60 = x
150+3x-75+2x-50-x+25+360 = 6x
50+4x+360 = 6x
410 = 2x
x = 205

Answer: 205$